The length of a rectangle is 5 ft more than twice the width, and the area of the rectangle is 63 ft^2Find the dimensions of the rectangle.
Length: _ft Width: _ft

Answer by checkley77(12844)

AREA=63
L=2W+5
AREA=LW
63=(2W+5)*W
63=2W^2+5W
2W^+5W-63=0
(2W-9)(W+7)=0
2W-9=0
W=9/2
W=4.5 ANS. FOR THE WIDTH.
L=2*4.5+5
L=9+5
L=14 ANS. FOR THE LENGTH.
PROOF:
63=14*4.5
63=63